OpenStudy (anonymous):

if the question asks "log .001" is that -3? is that a real solution?

OpenStudy (dinnertable):

Using the power rule of logarithms:\[\log(0.001) = \log(10^{-3}) = -3\log(10) = -3\]And yes this is a real solution because from a graphical perspective, the function exists for x > 0 which means that even though 0.001 is very small, it still exists.